Two small charged metal spheres are placed 20cm apart in air. If they have +20 x 10-6 C and +4 x 10-6 C charges on them respectively, what force acts on them? What is the electric field intensity halfway between them?


1
Expert's answer
2022-03-20T18:48:46-0400

The force between charges

"F=k\frac{Q_1Q_2}{r^2}=9*10^9*\frac{20*10^{-6}*4*10^{-6}}{0.2^2}=18\:\rm N"

The electric field intensity halfway between charges

"E=E_1-E_2=k\frac{Q_1}{r^2/4}-k\frac{Q_2}{r^2/4}"

"E=9*10^9*\frac{20*10^{-6}}{0.1^2}-9*10^9*\frac{4*10^{-6}}{0.1^2}\\
=1.44*10^7\:\rm N/C"